OpenStudy (anonymous):

sin a= 3/5 and is in quadrant II cos b = 2/5 and is in Quadrant I Find Cos (a-b)

OpenStudy (anonymous):

yo yo yo i still dont know how to do

OpenStudy (anonymous):

cos(a-b) = cos a cos b + sin a sin b

OpenStudy (anonymous):

sin b is positive (in first quad), cos a is neg (in second quad)

OpenStudy (anonymous):

use pythag to find co functions (if needed)

OpenStudy (anonymous):

wait so what is cos a and what is sina and sin b

OpenStudy (anonymous):

remember \[\cos ^{2}x+\sin ^{2}x=1\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

but what is it?

OpenStudy (anonymous):

sin a = 3/5, cos a < 0, so \[\cos a=-\sqrt{1-\left( \frac{ 3 }{ 5 } \right)^{2}}\]